# Imitation Learning with Limited Actions via Diffusion Planners and Deep Koopman Controllers

### v1

@@ III-C Koopman Operator Theory
More formally, let $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ denotes the state of a dynamical system at time $t$. Given a discrete-time dynamical system governed by $x_{t+1}=\mathbf{F}(x_{t})$, the Koopman Operator $\mathcal{K}:\mathcal{F}\rightarrow\mathcal{F}$ is defined as a linear transformation acting on observable functions $g$, where $\mathcal{K}g\triangleq g\circ\mathbf{F}$. Here $g:\mathcal{X}\rightarrow\mathbb{R}\in\mathcal{F}$ is the collection of all functions (also called observables) that forms an infinite-dimensional Hilbert Space. For any observable function $g$, there exists a Koopman Operator $\mathcal{K}$ such that,
@@
Koopman Operators have been extended to nonlinear control systems $x_{t+1}=\mathbf{F}(x_{t},a_{t})$ by linearly injecting the control input $a_{t}$ into the Koopman Observation space  [32, 33],
@@ IV Imitation Learning with Limited Action Data
Lifting the System and Latent Action Prediction. We use Deep Koopman embeddings to lift the system and assume that sequences of consecutive latent states encode action/control information. Let $x^{\tau}_{t}$ represent the trajectory around time $t$, consisting of past, current, and future states: $x^{\tau}_{t}=(x_{t-n},x_{t-n+1},\cdots,x_{t},x_{t+1},\cdots,x_{t+k})$. The state trajectory encoder (latent action predictor) $f_{\theta}$ maps these trajectories to control signals in the Koopman observation space:
$g_{\theta}(x_{t+1})$ $=\mathcal{K}_{\theta}g_{\theta}(x_{t})+f_{\theta}(x^{\tau}_{t})$ (3)
This linear structure encourages $f_{\theta}$ to learn the difference between lifted states, $g_{\theta}(x_{t+1})$ and $\mathcal{K}_{\theta}g_{\theta}(x_{t})$, without needing complex information bottlenecks.
We assume state-action trajectories are generated by the same system, so the observation function $g_{\theta}$ remains the same. We define an action decoder $d_{\phi}$:
@@
$\mathcal{L}_{kpm}=\frac{1}{T}\sum_{t=1}^{T}\|\mathcal{K}_{\theta}z_{x_{t}}+f_{\theta}(x_{t}^{\tau})-g_{\theta}{x_{t+1}}\|^{2}$ (6)
@@ V-A Benchmark Tasks
We use the Datasets with Diverse human Demonstrations for Imitation Learning (D3IL) benchmark dataset [5], which contains multi-modal human demonstrations and robot trajectories for various manipulation tasks. We used five tasks of varying difficulty:
  • Avoiding: The robot’s task is to move its end-effector from left to right across a table while avoiding fixed obstacles. The robot’s state is defined by the position of its end-effector, and its actions are based on controlling its end-effector velocity. Any collision with an obstacle leads to immediate failure.
  • Aligning: In this task, the robot must manipulate a block to match a specified target pose. The state consists of the robot’s end-effector position, the block’s pose, and the target pose, while actions control the end-effector’s velocity. The goal is to push or pull the block until it reaches the target configuration.
  • Sorting (4 blocks): The robot must sort four blocks (two red, two blue) by pushing them into their designated zones. The state includes the robot’s end-effector position and the pose of the blocks, with actions controlling the end-effector’s velocity. Failure occurs if a block is pushed into the wrong zone.
  • Stacking: This is a challenging task, requiring the robot to stack blocks in a specific order within a target zone. The state includes the end-effector position, gripper status (open or closed), block poses, and the required stacking order. Actions involve controlling the robot’s 7-DoF joint velocity (which is not observed) and the gripper. Placing blocks outside the target zone or in the wrong order results in failure.
  • Partially-Observable Sorting: A more difficult partially-observable version of original Sorting task, where the exact poses of the blocks are hidden and only their positions can be observed. In addition, Gaussian noise is added to robot’s end-effector positions in the state.
Due to variation in dataset size across tasks, for training, we use 50% of the action-labeled data for the smaller task datasets (avoiding and aligning) and 25% for larger datasets (sorting and stacking). We defined five different “levels” of action data based on dataset size. For the smaller tasks, the levels corresponded to (2%, 5%, 10%, 25%, 50%), while for the larger tasks, the levels were (1%, 2%, 5%, 10%, 25%).
@@ V-B Baselines
  • VAE [41]: A Variational Autoencoder method that trains a controller by encoding observational trajectories into an embedding. After training the VAE, an action decoder is trained to map sampled latent embeddings from the VAE to robot actions. This model can potentially learn useful action embeddings from the observations, but does not leverage dynamics.
@@
All methods, except for Diffusion Policy, follow a plan-then-control scheme. All plan-then-control methods use the same Decision Diffuser for a given task. For fair comparison, we use the same underlying model for the Diffusion Policy and Decision Diffuser, with the difference being that the Diffusion Policy predicts future actions instead of states. For KOAP, $g_{\theta}$ and $g_{\psi}$ were MLPs and $f_{\theta}$ was a LSTM. The decoder was $d_{\phi}$ was a linear model.
We use the original dataset provided by D3IL and sample fixed-length trajectory sequences. The first two states (no actions) are used as conditioning input for the planner and controller. The controllers are trained to predict a sequence of actions (12 steps). During inference, we employ receding-horizon control: the first four predicted actions are executed in the environment, with replanning and execution until the episode is completed. For complete implementation details, please refer to our online code repository [^1] To be made available after review..
@@ VI Case Study: Real Robot Scooping
We evaluate KOAP on a robot scooping task, where the goal is to use a spoon to manipulate a chocolate piece in a bowl. The experiment was conducted using a 7-DoF Panda robot.
Human demonstrations were collected by manually scooping the chocolate with the spoon, and robot action-labeled demonstrations were gathered through kinesthetic teaching using the same spoon. We collected 100 trajectories for both human and robot demonstrations. To reduce the computational overhead of generating pixel states for planning, we use a pretrained ResNet[43] to extract feature embeddings from pixel observations. The diffusion planner generates future embeddings conditioned on the current pixel observations, and the controller produces robot joint angles as actions.

### v2

@@ III-C Koopman Operator Theory
More formally, let $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ denotes the state of a dynamical system at time $t$. Given a discrete-time dynamical system governed by $x_{t+1}=\mathbf{F}(x_{t})$, let $\mathcal{F}$ denote the collection of observable functions $g:\mathcal{X}\rightarrow\mathbb{R}$ that forms an infinite-dimensional Hilbert space. The Koopman Operator $\mathcal{K}:\mathcal{F}\rightarrow\mathcal{F}$ is defined as a linear transformation acting on these observables, such that $\mathcal{K}g\triangleq g\circ\mathbf{F}$ for any $g\in\mathcal{F}$. The Koopman Operator transforms observable functions according to the system dynamics, such that:
@@
Koopman Operators have been extended to nonlinear control systems
$x_{t+1}=\mathbf{F}(x_{t},a_{t})$ by linearly injecting the control input $a_{t}$ into the Koopman Observation space  [32, 33, 34],
@@ IV Imitation Learning with Limited Action Data
Lifting the System and Latent Action Prediction. We use Deep Koopman embeddings to lift the system and assume that sequences of consecutive latent states encode action/control information. Let $x^{\tau}_{t}$ represent a trajectory around time $t$, consisting of past, current, and a near future state $x_{t+k}$: $x^{\tau}_{t}=(x_{t-n},x_{t-n+1},\cdots,x_{t},x_{t+k})$. The state trajectory encoder (latent action predictor) $f_{\theta}$ maps such trajectories to control signals in the Koopman observation space:
$g_{\theta}(x_{t+1})$ $=\mathcal{K}_{\theta}\times g_{\theta}(x_{t})+f_{\theta}(x^{\tau}_{t})$ (3)
The design of state trajectory allows near future state to provide information about the action to take, but disallows direct feeding in the future state $x_{t+1}$, which prevents learning trivial representation that predicts $x_{t+1}$ directly. The linear structure encourages $f_{\theta}$ to learn the difference between lifted states, $g_{\theta}(x_{t+1})$ and $\mathcal{K}_{\theta}\times g_{\theta}(x_{t})$, without needing complex information bottlenecks. We assume state-action trajectories are generated by the same system, so the observation function $g_{\theta}$ remains the same. We define an action decoder $d_{\phi}$:
@@
$\mathcal{L}_{kpm}=\frac{1}{T}\sum_{t=1}^{T}\|\mathcal{K}_{\theta}\times g_{\theta}({x_{t}})+f_{\theta}(x_{t}^{\tau})-g_{\theta}({x_{t+1}})\|^{2}$ (6)
@@ V-A Benchmark Tasks
We use the Datasets with Diverse human Demonstrations for Imitation Learning (D3IL) benchmark [5], containing multi-modal human demonstrations and robot trajectories across five manipulation tasks of varying difficulty:
  • Avoiding: Robot navigates from left to right while avoiding fixed obstacles. State includes end-effector position; actions control end-effector velocity. Collisions result in failure.
  • Aligning: Robot manipulates (push or pull) a block to match a target pose. State includes end-effector position, block pose, and target pose; actions control end-effector velocity.
  • Sorting (4 blocks): Robot sorts four blocks (two red, two blue) into designated zones. State includes end-effector position and block poses; actions control end-effector velocity. Placing blocks in wrong zones causes failure.
  • Stacking: This is a difficult task. Robot stacks blocks in a specific order within a target zone. State includes end-effector position, gripper status, block poses, and stacking order; actions control 7-DoF joint velocity (which is not observed in state) and gripper. Placing blocks outside the target zone or in the wrong order results in failure.
  • Partially-Observable Sorting: A more difficult partially-observable version of original Sorting task, where only block positions (not poses) are observable, with added Gaussian noise to end-effector positions in the state.
Due to big variation in dataset size across tasks, for training, we use 50% of the action-labeled data for the smaller task datasets (avoiding and aligning) and 25% for larger datasets (sorting and stacking). We defined five different “levels” of action data based on dataset size. For the smaller tasks, the levels corresponded to (2%, 5%, 10%, 25%, 50%), while for the larger tasks, the levels were (1%, 2%, 5%, 10%, 25%).
@@ V-B Baselines
  • VAE [42]: A Variational Autoencoder method that trains a controller by encoding full observational trajectories into an embedding. After training the VAE, an action decoder is trained to map sampled latent embeddings from the VAE to robot actions. This model can potentially learn useful action embeddings from the observations, but does not leverage dynamics.
@@
All methods, except for Diffusion Policy, follow a plan-then-control scheme. All plan-then-control methods use the same Decision Diffuser for a given task. For fair comparison, we use the same underlying model for the Diffusion Policy and Decision Diffuser, with the difference being that the Diffusion Policy predicts future actions instead of states. For KOAP, $g_{\theta}$ and $g_{\psi}$ were MLPs and $f_{\theta}$ was a LSTM. The decoder $d_{\phi}$ was a linear model.
We use the original dataset provided by D3IL and sample fixed-length trajectory sequences. The first two states (no actions) are used as conditioning input for the planner and controller. For future state $x_{t+k}$ in state trajectory, we find $k=8$ performs well empirically. For fair comparison, all controllers are trained to predict action chunks (with 12 steps). During inference, we employ receding-horizon control: the first 4 predicted actions are executed in the environment, with replanning and execution until the episode is completed. For complete details on the implementation, see our online code repository.
@@ VI Case Study: Real Robot Scooping
We evaluate KOAP on a robot scooping task, where the goal is to use a spoon to manipulate a chocolate piece in a bowl. The experiment was conducted using a 7-DoF Panda robot. Human demonstrations were collected by manually scooping the chocolate with the spoon, and robot action-labeled demonstrations were gathered through kinesthetic teaching using the same spoon. We collected 100 trajectories for both human and robot demonstrations. To reduce the computational overhead of generating pixel states for planning, we use a pretrained ResNet [44] to extract feature embeddings from pixel observations. The diffusion planner generates future embeddings conditioned on the current pixel observations, and the controller produces robot joint states as actions.
@@ VII Conclusion
ACKNOWLEDGMENT
This research is supported by the National Research Foundation, Singapore under its Medium Sized Center for Advanced Robotics Technology Innovation.